OpenStudy (anonymous):

find f'(x) for f(x)=sqrt(4+e^2x)

OpenStudy (anonymous):

f(x) = u^1/2 f'(x) = (1/2) [u^-1/2](du/dx) chain rule du/dx = 2e^2x